A plant is going to buy a new machine. A productivity analysis is being done for the purchase.  

The current machine produces 120 units per day. It needs 40 labor hrs per day at a labor rate of $20 per hr and $500 capital cost per day.

After automation, it is estimated to take 16 labor hrs and $1400 capital cost per day. It will then produce 200 units per day. Labor rate increases to $25/Hr.

Consider only labor and capital inputs. Ignore all other inputs. Find the labor, capital and total factor productivity now and after automation. Would you recommend the automation project? Why and why not?
Answer #1

OLD PRODUCTIVITY

VALUES
LABOR = 800(40 * 20)

CAPITAL = 500


OUTPUT = 120

PRODUCTIVITY = OUTPUT / TOTAL INPUT

PRODUCTIVITY = 120 / (800 + 500) = 0.0923


NEW PRODUCTIVITY

VALUES
LABOR = 400(16 * 25)

CAPITAL = 1400


OUTPUT = 200

PRODUCTIVITY = OUTPUT / TOTAL INPUT

PRODUCTIVITY = 200 / (400 + 1400) = 0.1111


CHANGE IN PRODUCTIVITY

CHANGE = ((NEW - OLD) / OLD) * 100

% = ((0.1111 - 0.0923) / 0.0923 = 20.37


SINCE THERE IS A 20.37% INCREASE IN PRODUCTIVITY, WE SHOULD INVEST IN THE NEW METHOD TO INCREASE PRODUCTIVITY.